Tropical Cable and Conductors Ltd can sell x items per day at a price of p GH¢ each, where  

p =125 – 5x/3 . The cost of production for x items is 500 + 13x +0.2x2

1.    Find the marginal cost function and the total revenue function

2.    Find how much they should be produced to have a maximum profit.

3.    Assuming that all items produced can be sold. What is the maximum profit?

4.    How much should be produced to breakeven?

5.    Calculate the price that maximizes profit.

6.    Compute the price elasticity of demand for the firm’s product at the profit maximizing output level.

7.    As a consultant advise the firm on its pricing policy.

8.    The government imposes a per unit tax on the firm, calculate:

a. the new profit maximizing level of output.

b. the new maximum profit.



1
Expert's answer
2021-11-07T19:44:09-0500

1. "MC=\frac{\Delta TC}{\Delta Q}=13+0.4x"

"TR=P\times Q\\=x({125-\frac{5x}{3}})\\=125x-\frac{5x^2}{3}"